# Gaussian Herding across Pens:::: An Optimal Transport Perspective on Global Gaussian Reduction for 3DGS

### v1

@@ 3.3 Training Details with GHAP Algorithm
Integrating these components, we present our complete training pipeline in Algorithm 2, called Gaussian Herding Across Pens (GHAP). The process begins with standard 3DGS optimization for $T$ steps, followed by blockwise GMR. We then freeze the geometric parameters ($\mu$ and $\Sigma$) while fine-tuning appearance attributes (opacity $\alpha$ and color $c$) through an additional $T$-step 3DGS optimization. This procedure can be iteratively applied throughout training as needed.
@@ 4.1 Experimental Setup
Datasets
For a comprehensive evaluation of our GHAP algorithm, we use three real-world datasets: Tanks and Temples [29], Mip-NeRF 360 [30], and Deep Blending [31], which cover varying levels of detail, lighting conditions, and scene complexities. For each dataset, we adopt the same scenes as in [8].
  • Tanks and Temples: We evaluate two unbounded outdoor scenes, “Truck” and “Train”, both featuring centered viewpoints.
@@
Baselines
To evaluate the effectiveness of our proposed method, we compare it against several strong baselines in the domain of 3DGS compaction: Mini-Splatting-D [12], AtomGS [13], LightGaussian [11], and PUP-3DGS [15]. We set $s=3\times 10^{3}$. The sensitivity analysis of $s$ is given in Appendix. The initialization of the iterative algorithm in (1) randomly selects $m$ Gaussians per block.
Evaluation Metrics We assess 3DGS compaction using standard metrics for rendering quality and efficiency. For quality, we report: (1) PSNR, measuring pixel-level accuracy; (2) SSIM, evaluating perceptual similarity based on luminance, contrast, and structure; and (3) LPIPS, capturing perceptual distance via a learned model. Higher PSNR/SSIM and lower LPIPS indicate better quality. For efficiency, we measure FPS across retention ratios, where lower retention yielding higher FPS implies greater rendering efficiency.
Device
All experiments were conducted on a server with 256 GB RAM and a 96-core Intel® Xeon® Platinum 8255C CPU, and on a workstation equipped with five NVIDIA RTX 3090 GPUs, each with 24 GB of VRAM.
@@ 4.2 Quantitative Results
We first demonstrate the effectiveness of our GHAP compaction method as a plug-in module within 3DGS and its variants. Second, we show that our approach outperforms other SOTA compaction techniques. We present the results first and summarize our findings at the end of the section.
As a Plug-In Compaction Method
Our method can be used as a plug-in compaction method in various 3DGS training algorithms. This demonstrates the broad applicability of our proposed method. To verify this, we apply our compaction method within various 3DGS pipelines. Specifically, we consider three representative variants: 3DGS-30K, AtomGS, and Mini-Splatting-D. Each of them employs a distinct densification strategy, and our method can be directly embedded into the pipeline without extensive engineering effort.
For each variant, we evaluate performance under two retention ratios (10% and 20%) and report results using three standard metrics: PSNR, SSIM, and LPIPS. Table 1 summarizes our experimental results.
Table 1: Quantitative evaluation of our GHAP method with respect to compaction performance applied to different 3DGS variants under varying retention ratios (10% and 20%).
Method $\rho$ Tanks&Temples MipNeRF-360 Deep Blendering
@@
3DGS-30k 10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
Mini-Splatting-D 10% 0.835 23.232 0.198 0.802 27.090 0.250 0.909 30.042 0.254
20% 0.855 23.403 0.171 0.821 27.310 0.214 0.912 30.170 0.238
100% 0.848 23.338 0.140 0.832 27.486 0.176 0.907 29.980 0.211
AtomGS 10% 0.793 22.988 0.274 0.764 26.535 0.307 0.899 29.347 0.282
20% 0.812 23.282 0.240 0.788 27.025 0.269 0.902 29.347 0.268
100% 0.814 23.289 0.235 0.796 27.135 0.251 0.902 29.314 0.267
Comparison with SOTA Compaction Methods
We compare our method with two recent compaction baselines: PUP-3DGS and LightGaussian. To ensure a fair evaluation, all methods are built upon the same backbone model, 3DGS-30K. Compaction is performed at the same stage, when the number of Gaussians reaches 15K, followed by 15K fine-tuning iterations. During fine-tuning, we use identical optimization settings across all methods, including learning rate and other hyperparameters. Two compaction ratios are considered: 10% and 20%. A detailed comparison of rendering quality across different compaction methods and retention ratios is presented in Table 2.
Table 2: We compare our GHAP method against LightGaussian and PUP-3DGS, modifying their original workflows to align with our compaction pipeline.
Method-Compact Rate Tandt&Temples MipNeRF-360 Deep Blendering
@@
3DGS-30k-100% 0.853 23.785 0.169 0.813 27.554 0.221 0.907 29.816 0.238
GHAP(Ours)-10% 0.818 23.312 0.242 0.764 26.404 0.314 0.905 29.647 0.264
Light-GS-10% 0.791 22.560 0.248 0.779 26.585 0.261 0.861 27.136 0.326
Pup-GS-10% 0.767 21.519 0.28 0.753 25.332 0.309 0.895 29.153 0.274
GHAP(Ours)-20% 0.835 23.615 0.212 0.788 26.973 0.275 0.907 29.864 0.252
Light-GS-20% 0.791 22.559 0.248 0.779 26.588 0.261 0.861 27.142 0.326
Pup-GS-20% 0.809 22.603 0.228 0.790 26.671 0.257 0.905 29.719 0.248
Figure 4: Visualization before and after compaction. Various scenes before and after compaction under different 3DGS algorithms. Our GHAP method yields negligible loss in rendering quality.
Takeaways
Based on these experiments, we draw four key observations:
  • Superior fidelity retention. Our method substantially outperforms both LightGaussian and PUP-3DGS in SSIM and PSNR. Surprisingly, our method can even achieve better performance than the original method, this is demonstrated in Table 1 on the Deep Blending scene with the Mini-Splatting-D backbone and at a 10% retention ratio. Likewise, in Table 2, we consistently surpass LightGaussian and PUP-3DGS in most datasets and metrics.
  • Efficiency gains. At a 10% Gaussian primitives retention rate, our compaction framework yields a 140.3% increase in rendering speed (FPS), demonstrating that significant model slimming can be achieved without sacrificing and even often improving performance (see ablation study).
  • Cross-dataset consistency. Performance remains stable across datasets. In outdoor scenes (Tanks and Temples), SSIM value fidelity drops ratio (relative drop in rendering metrics after compaction) less than 4.1%. Meanwhile, for the indoor scene, the PSNR value retains 99.4% of the original value.
  • LPIPS sensitivity. We observe that our GMR compaction method incurs a relatively larger degradation in LPIPS compared to SSIM and PSNR, although the reduction remains within acceptable bounds. This trend aligns with prior findings [11, 15], as LPIPS is more sensitive to structural and textural distortions. Since compaction inherently reduces model capacity, fine-grained texture reconstruction becomes more challenging, which explains the behavior observed in our qualitative and quantitative analyses
Qualitative Visualization
@@
We visualize two scenes for each baseline 3DGS algorithm before and after compaction in Figure 4. Across most scenes, our compaction method successfully preserves rendering quality using only 10% of the Gaussian primitives. By design, our reduced Gaussians approximate the original Gaussians from the 3DGS algorithm. If the original model lacks good visual quality due to lack to Gaussian primitives in certain regions, our method cannot perform well—as demonstrated in the ‘Bicycle’ scene, where the poor performance of the Atom-GS baseline prevents our approach from maintaining high fidelity after compaction. Interestingly, in some cases, our post-compaction results surpass the pre-compaction quality. A representative example is the ‘Kitchen’ scene based on Mini-Splatting-D. We conjecture this occurs because Mini-Splatting-D lacks a pruning mechanism, often generating an oversaturated set of Gaussians that introduce visual artifacts (e.g., unnatural shadows in the lower-left region). Our method mitigates this issue by globally regularizing the Gaussian distribution: improving expressiveness, reducing unnecessary density, and preserving the underlying 3D structure.
@@ 4.3 Ablation Studies
GMR versus Random Subsampling We evaluate the contribution of each stage in our pipeline—geometric compaction and appearance optimization—on the Tanks and Temples dataset. For comparison, we use random subsampling at the same compaction ratio. This allows us to assess the effectiveness of our design choices at each stage. The experiments show that our compaction method is significantly more effective than random subsampling,
@@
producing a more meaningful scene representation. Appearance optimization further leads to substantial quantitative improvements. Together, these results validate the necessity and effectiveness of both stages in our proposed pipeline.
Figure 5: GHAP efficiency-quality trade-off. SSIM and FPS (top) and LPIPS and PSNR (bottom) across different retention ratios $\rho$.
Rendering Metrics under GHAP
To analyze the efficiency-quality trade-off of the proposed GHAP approach, we vary the retention ratio $\rho$ from $10\%$ to $100\%$. A lower retention ratio enables more aggressive compaction, resulting in faster rendering speed but potentially compromising rendering quality. Our results demonstrate that decreasing the retention ratio yields 2×2$\times$2 × faster rendering with only marginal quality degradation. For instance, as $\rho$ decreases from $100\%$ to $10\%$: LPIPS increases modestly from 0.14 to 0.21; SSIM declines slightly from 0.885 to 0.860; PSNR remains stable between 25.0 dB and 25.5 dB. These findings indicate that retaining even a modest number of Gaussians effectively preserves visual quality, achieving an excellent performance-fidelity trade-off.
5 Conclusion
We present a novel optimal transport-based global Gaussian reduction approach for 3D Gaussian Splatting, addressing the inefficiencies of existing Gaussian pruning and compaction methods. By formulating the problem as Gaussian mixture reduction, our method minimizes composite transport divergence to derive a compact geometric representation, followed by appearance fine-tuning with significantly fewer primitives. It provides a theoretically sound foundation for efficient neural rendering. Extensive experiments demonstrate that our approach achieves rendering quality nearly indistinguishable from vanilla 3DGS while retaining only  10% of the Gaussians, outperforming existing compaction techniques. Numerical results show our framework is versatile, compatible with various 3DGS pipelines. This work advances the practical deployment of 3DGS by striking an optimal balance between computational efficiency and visual fidelity. Future work includes extending our method to dynamic 3DGS for real-time temporal scenes and developing adaptive reduction strategies for task-specific needs.

### v2

@@ 3.3 Training Details with GHAP Algorithm
Integrating these components, we present our complete training pipeline in Algorithm 2, called Gaussian Herding Across Pens (GHAP). The process begins with standard 3DGS optimization for $T$ steps, followed by blockwise GMR. We then freeze the geometric parameters ($\mu$ and $\Sigma$) while fine-tuning the appearance attributes (opacity $\alpha$ and color $c$) through an additional $T$-step 3DGS optimization. This procedure can be applied iteratively throughout training as needed.
@@ 4.1 Experimental Setup
Datasets. For a comprehensive evaluation of GHAP algorithm, we use three real-world datasets: Tanks & Temples (knapitsch2017Tanks, ), Mip-NeRF 360 (hedman2018deep, ), and Deep Blending (barron2022mip, ), which cover varying levels of detail, lighting conditions, and scene complexities. For each dataset, we adopt the same scenes as in (bagdasarian20243dgs, ).
  • Tanks & Temples: We evaluate two unbounded outdoor scenes, “Truck” and “Train”, both featuring centered viewpoints.
@@
Baselines. To evaluate the effectiveness of our proposed method, we compared it against four strong compaction techniques: LightGaussian (fan2024lightgaussian, ), PUP-3DGS (hanson2024pup, ), Trimming the Fat (ali2024trimming, ), and MesonGS (xie2024mesongs, ), as well as four end-to-end 3DGS variants: Mini-Splatting(-D) (fang2024mini, ), AtomGS (liu2024atomgs, ), 3DGS-MCMC (kheradmand20243d, ), and LocoGS (shin2025locality, ). Notably, Mini-Splatting-D and AtomGS were employed as backbone models in our approach, while the other variants were used for direct comparative evaluation against the compaction methods. A consistent evaluation protocol was established to ensure fair and reliable conclusions. For post-training compaction methods applicable to pre-trained models–including GHAP, LightGaussian, PUP-3DGS, Trimming the Fat, and MesonGS–we initialized all from the same backbone model (trained using vanilla 3DGS, Mini-Splatting-D, or AtomGS for 15k iterations) and applied their respective compaction procedures directly, excluding any compression-specific modules. All models subsequently underwent identical fine-tuning for 15k iterations to achieve the target retention ratio. For the other end-to-end variants (MiniSplatting, 3DGS-MCMC and LocoGS), we executed training for 30k iterations under their default configurations. Detailed experimental steps for each method can be found in the appendix.
Evaluation Metrics. We assess 3DGS compaction using standard metrics for rendering quality. We report: (1) PSNR, measuring pixel-level accuracy; (2) SSIM, evaluating perceptual similarity based on luminance, contrast, and structure; and (3) LPIPS, capturing perceptual distance via a learned model. Higher PSNR/SSIM and lower LPIPS indicate better quality. For each method, we also report the corresponding number of Gaussian primitives.
@@ 4.2 Quantitative Results
We first show that our approach outperforms other SOTA compaction techniques. Second, we demonstrate the effectiveness of our GHAP compaction method as a plug-in module within 3DGS and its variants. We present the experimental results followed by our key findings.
Comparison with SOTA. We compare our method against a comprehensive set of baselines, which can be categorized into two groups:
Our approach is evaluated in two configurations: 3DGS+GHAP and MiniSplatting+GHAP. The former uses the same vanilla 3DGS backbone and 10% retention rate as the pruning baselines for a direct comparison of compaction strategies. The latter replaces the built-in pruning step in Mini-Splatting with our GHAP algorithm to demonstrate its effectiveness on a different backbone.
Quantitative results are summarized in Table LABEL:tab:sota_comparison, with methods grouped by their final primitive count. As expected, LocoGS achieves strong performance due to its larger primitive count. Among methods with comparable primitive counts (Group 2), our 3DGS+GHAP achieves superior performance in SSIM and PSNR, with a marginally lower LPIPS score. The advantage of our compaction strategy is further evident when using the Mini-Splatting backbone. Our MiniSplatting+GHAP outperforms other compaction-based approaches while often using fewer primitives. As shown in Fig. 4 (left), this performance lead is consistent across a wide range of retention ratios, not just at $\rho=0.1$.
Table 1: Quantitative comparison of rendering quality against SOTA methods. Our method (GHAP), applied to two different backbones (3DGS and Mini-Splatting), consistently matches or surpasses pruning-based baselines, even at a low retention rate (10%), while maintaining competitive runtime and memory usage. The best results for a given number of Gaussians are shown in bold; second best are underlined.
Method Tanks&Temples MipNeRF-360 Deep Blending
@@
Vanilla 3DGS 0.853 23.785 0.169 1577 0.813 27.554 0.221 2627 0.907 29.816 0.238 2475
LocoGS 0.843 23.655 0.191 571 0.798 27.049 0.257 674 0.903 29.972 0.261 529
3DGS+GHAP (ours) 0.818 23.312 0.242 157 0.764 26.404 0.314 263 0.905 29.647 0.264 248
LightGaussian 0.756 22.113 0.306 158 0.735 25.674 0.331 263 0.869 28.010 0.327 248
PUP-3DGS 0.767 21.519 0.280 158 0.753 25.332 0.309 262 0.895 29.153 0.274 248
Trimming the Fat 0.776 21.535 0.293 156 0.731 25.255 0.343 263 0.887 28.056 0.302 247
MesonGS 0.811 20.714 0.208 157 0.773 24.924 0.264 263 0.896 28.693 0.264 248
3DGS-MCMC 0.779 22.141 0.282 157 0.763 25.957 0.309 263 0.885 28.976 0.298 248
MiniSplatting 0.799 22.661 0.265 78 0.759 26.022 0.318 111 0.895 29.395 0.289 125
MiniSplatting+GHAP (ours) 0.835 23.232 0.198 79 0.802 27.090 0.250 112 0.909 30.042 0.254 127
Figure 4: Comparison of compaction methods on the Tanks & Temples dataset. Left: Rate-Distortion (RD) curves; middle: computational time, and right: memory consumption.
Runtime & Memory Usage Comparison. Crucially, the improved performance of our method does not come at a computational cost. As depicted in Fig. 4 (middle), our method’s runtime is faster than all baselines except the exceptionally swift Trimming the Fat. While our method exhibits a slightly higher memory footprint (in Fig. 4 right) during compaction due to pairwise distance computation in each KD-tree block, the difference is not substantial (less than an order of magnitude).
As a Plug-In Compaction Method. Our method can be used as a plug-in compaction method in various 3DGS training algorithms. This demonstrates the broad applicability of our proposed method. To verify this, we apply our compaction method within various 3DGS pipelines. Specifically, we consider three representative variants as the backbone: 3DGS (kerbl20233d, ) , AtomGS (liu2024atomgs, ), and Mini-Splatting-D (fang2024mini, ). Each of them employs a distinct densification strategy, and our method can be directly embedded into the pipeline without extensive engineering effort. For each backbone, we evaluate performance under two retention ratios (10% and 20%). Tab. 2 summarizes our experimental results.
Table 2: Quantitative results with different backbones. The compaction performance of our GHAP method when used with different 3DGS variants as backbones under varying retention ratios.
Backbone $\rho$ Tanks&Temples MipNeRF-360 Deep Blending
@@
3DGS-30k 10% 0.818 23.312 0.242 157 0.764 26.404 0.314 263 0.905 29.647 0.264 248
20% 0.835 23.615 0.212 314 0.788 26.973 0.275 526 0.907 29.864 0.252 596
100% 0.853 23.785 0.169 1577 0.813 27.554 0.221 2627 0.907 29.816 0.238 2475
Mini-Splatting-D 10% 0.835 23.232 0.198 313 0.802 27.090 0.250 357 0.909 30.042 0.254 331
20% 0.855 23.403 0.171 626 0.821 27.310 0.214 614 0.912 30.170 0.238 662
100% 0.848 23.338 0.140 3132 0.832 27.486 0.176 3578 0.907 29.980 0.211 3316
AtomGS 10% 0.793 22.988 0.274 189 0.764 26.535 0.307 293 0.899 29.347 0.282 271
20% 0.812 23.282 0.240 378 0.788 27.025 0.269 586 0.902 29.347 0.268 542
100% 0.814 23.289 0.235 1897 0.796 27.135 0.251 2928 0.902 29.314 0.267 2709
The quantitative results in Tab. 2 demonstrate that our method effectively preserves the backbone models’ visual quality, even at an extreme retention rate of 10%. Notably, on some scenes (highlighted in bold), the compacted model’s performance surpasses that of the uncompacted backbone.
Beyond quantitative metrics, we provide a qualitative analysis by visualizing multiple scenes before and after compaction in Fig. 5.
Figure 5: Visual Quality Before and After Compaction. Visual results for various scenes under different 3DGS backbones, compacted to 10% of their primitives using our GHAP method. Our approach preserves rendering quality with negligible loss. In some cases (e.g., “Kitchen”), compaction even improves quality by regularizing an over saturated Gaussian distribution.
@@
As evidenced in the figure, our method successfully preserves rendering quality across most scenes while using only 10% of the Gaussian primitives. Interestingly, in certain cases, our compaction not only preserves but surpasses the original quality. A representative example is the “Kitchen” scene (Mini-Splatting-D backbone). We conjecture this improvement occurs because Mini-Splatting-D lacks a pruning mechanism, often generating an over saturated set of Gaussians that introduce visual artifacts (e.g., the unnatural shadows in the lower-left region). Our method acts as a global regularizer, mitigating this issue by reducing unnecessary density while improving the overall expressiveness and preserving the underlying 3D structure. Naturally, our approach is inherently limited by the quality of its input. If the original model suffers from significant artifacts due to a lack of primitives in certain regions, our compaction cannot resolve these fundamental issues. This limitation is demonstrated in the “Bicycle” scene, where artifacts present in the Atom-GS backbone persist after compaction.
@@ 4.3 Ablation Studies
Influence of KD Tree Depth. To validate the effectiveness of the KD-tree partitioning strategy, we provide an ablation study on it. Due to the large scale of the three primary datasets, low KD-tree depths result in an excessive number of points per block, making it infeasible to run the GMR algorithm. Therefore, we conduct ablation experiments on the smaller mic scene from NeRF-Synthetic. Results in Fig. 6 show that increasing KD-tree depth reduces memory usage and runtime, while PSNR first improves and then declines. This indicates that moderately finer partitions allow GMR to compact regions more effectively, whereas overly fine splits may fragment primitives and degrade quality.
Figure 6: Hyperparameters. The impact of KD tree depth on runtime performance, memory footprint, and rendering quality.
Loss Function Design. As shown in Tables LABEL:tab:sota_comparison and 2, our method exhibits a slight underperformance on the LPIPS metric. To address this, we investigate whether incorporating an LPIPS loss term can yield improvements.
Our baseline loss function follows the vanilla 3DGS formulation, using an L1-to-SSIM ratio of 8:2. We experiment with two new weighting schemes that include LPIPS: (1) L1:SSIM:LPIPS = 8:1:1, and (2) L1:SSIM:LPIPS = 6:2:2. The results of this ablation study on the Tanks&Temples dataset are shown in Fig. 7 (with additional results in the Appendix). Our analysis reveals a trade-off between perceptual and distortion metrics: increasing the weight of the LPIPS loss improves LPIPS scores but leads to a slight deduction in PSNR and SSIM.
Figure 7: Comparison of different loss weights. A trade-off between perceptual and distortion metrics.
GMR versus Random Subsampling. We conduct an ablation study on the Tanks&Temples dataset to evaluate the contribution of each stage in our pipeline: geometric compaction and appearance optimization. A random subsampling baseline is included to assess the effectiveness of our design choices.
@@
The results, presented in Table 3, compare two compaction schemes followed by the same fine-tuning procedure. Our findings demonstrate that: first, our compaction procedure is significantly more effective than random subsampling and other pruning baselines (as shown in Tab. LABEL:tab:sota_comparison). Second, the subsequent appearance optimization stage provides substantial quantitative improvements for both compaction approaches. The significant performance gain over the baseline validates the necessity and effectiveness of both stages in our proposed pipeline.
5 Conclusion and Discussion
We propose an optimal transport-based Gaussian mixture reduction framework for 3D Gaussian Splatting, achieving compact yet faithful representations. By minimizing composite transport divergence with appearance fine-tuning, our method preserves high visual fidelity while retaining only  10% of Gaussians, outperforming prior compaction techniques. The framework scales efficiently via block-wise KD-tree partitioning and integrates seamlessly with diverse 3DGS pipelines.
Future directions include enhancing robustness across challenging scene types, incorporating perceptual objectives, developing multi-scale and overlap-aware partitioning, adopting auto-tuned schedules, and extending to dynamic 3DGS for real-time temporal rendering.